OpenStudy (anonymous):

Characterization of exponential functions. I just don't understand this concept.

OpenStudy (anonymous):

Characterization of exponential functions. If for equal increments of a of the argument x, the values of the function f(x) change by the same ratio, that is f(x + a)/f(x) = b^(a) is constant (independent of x, but dependent on a), then the function is exponential with: P = f(0) as initial value a = f(1)/f(0) as base f(x) = Pa^(x) ^ I do not understand how they came up with f(x +a)/f(x) = b^(a)

OpenStudy (anonymous):

This is the concept above.

OpenStudy (anonymous):

sort of odd way to put it

OpenStudy (anonymous):

saying that your increase/ decrease is proportional

OpenStudy (anonymous):

if you think about it backwards it makes sense. suppose you have an exponential function that looks like say \[f(x)=50\times (.85)^{\frac{x}{4}}\] then \[f(x+a)=50\times (.85)^{\frac{x+a}{4}}=50\times (.85)^{\frac{x}{4}}\times (.85)^{\frac{a}{4}}\] so \[\frac{f(x+a)}{f(x)}=(.85)^{\frac{a}{4}}\]

OpenStudy (anonymous):

a constant, independent of x

OpenStudy (anonymous):

So what does this mean, in english?

OpenStudy (anonymous):

How come you used xz for the variable in the exponent, when the function is f(x)